OpenStudy (anonymous):

sin(x+y)sin(x-y)=2sinxcosy

zepdrix (zepdrix):

Frankyyyyy, eyyyyy. What do we do? Prove this identity?

OpenStudy (anonymous):

yes sir

zepdrix (zepdrix):

I guess we could apply our `Angle Sum Formula` and `Angle Difference Formula` for sine, and expand out all the stuff we get. Shouldn't be too bad, just a lot of multiplying.

OpenStudy (anonymous):

i expanded it to sinxcosy-cosxsiny+sinxcosy-cosxsiny.

OpenStudy (anonymous):

oh ok

zepdrix (zepdrix):

So that's what you did.. ok ok interesting.

OpenStudy (anonymous):

i just followed the sum and difference formulas

zepdrix (zepdrix):

You messed up a sign somewhere, hmm.

zepdrix (zepdrix):

\[\rm \color{red}{\sin(x+y)}\color{royalblue}{\sin(x-y)}\]So our formulas give us,\[\rm \color{red}{[\sin x \cos y+\sin y \cos x]}\color{royalblue}{[\sin x \cos y- \sin y \cos x]}\]

zepdrix (zepdrix):

Oh oh oh, these are being added together right?\[\large\rm \sin(x+y)+\sin(x-y)\]Otherwise we're getting a big huge mess I think. Was that just a typo?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\rm \color{red}{\sin(x+y)}+\color{royalblue}{\sin(x-y)}\]So then our expression becomes,\[\rm \color{red}{\sin x \cos y+\sin y \cos x}+\color{royalblue}{\sin x \cos y- \sin y \cos x}\]

zepdrix (zepdrix):

Ya I think you had the wrong sign on the second term, should be positive, ya?

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

Understand what the next step is?

OpenStudy (anonymous):

no thats where im stuck

zepdrix (zepdrix):

We actually have some like-terms, even though it might not seem like it. Notice that we have a siny*cosx and a -siny*cosx, ya?

OpenStudy (anonymous):

oh yeah they just cancel out

zepdrix (zepdrix):

Good good good. And you're left with sinx*cosy + sinx*cosy right? :)

OpenStudy (anonymous):

right

zepdrix (zepdrix):

again, combining like-terms, just two of those. yay team!

OpenStudy (anonymous):

wow. i thought too much of it. thanks.

OpenStudy (anonymous):

can we do just 1 more?

OpenStudy (anonymous):

tan(x+pi) - tan(pi-x) = 2tanx

zepdrix (zepdrix):

Do they want us to use our Angle Sum/Difference formulas again? The tangent ones are kind of annoying. We can do something different in this case if you're comfortable with it. Tangent is `periodic in pi`. Which means that if you add pi to an angle, you get the same result as before,\[\large\rm \tan(x+\pi)=\tan x\]

OpenStudy (anonymous):

my teacher wants me too, but I too think it's very annoying.

zepdrix (zepdrix):

Ughhh D:

OpenStudy (anonymous):

lets just do what you want to do though. For times sake

zepdrix (zepdrix):

Lemme finish explaining it this way first, it'll take like 1 minute, and then we can move onto the longer approach if you like. :)

zepdrix (zepdrix):

\[\large\rm \tan\left[\pi-x\right]=\tan\left[-(x-\pi)\right]\]Tangent is an `odd function`, so it follows this property: \(\rm tan(-a)=-tan(a)\) the negative can come outside.\[\large\rm \tan\left[-(x-\pi)\right]=-\tan(x-\pi)\]And again, since tangent is periodic in pi, x-pi is the same as x,\[\large\rm =-\tan(x)\]

zepdrix (zepdrix):

\[\large\rm \color{orangered}{\tan(x+\pi)}-\color{royalblue}{\tan(\pi-x)}\quad=\quad \color{orangered}{\tan x}-\color{royalblue}{(-\tan x)}\]

zepdrix (zepdrix):

=2tanx. Ok so maybe that way is a little more complicated than I was letting on :)

OpenStudy (anonymous):

i find this much better.

zepdrix (zepdrix):

Ya as long as your teacher is ok with it :) hehe

OpenStudy (anonymous):

im sure she doesnt mind seeing different ways that this problem can be solved. Thanks a ton!

zepdrix (zepdrix):

np